OpenStudy (anonymous):

tanA=x , sec2A = ? a. (1 + x^2)/(1 - x^2) b. (1-x^2)/2x

OpenStudy (anonymous):

a)

OpenStudy (anonymous):

how did u get that?? pls tell

OpenStudy (anonymous):

Sec2A=(1+tan^2A)/(1-tan^2A) so Sec2A=1 + x^2)/(1 - x^2)

OpenStudy (anonymous):

okie...thank u.. :)

OpenStudy (anonymous):

My pleasure........

OpenStudy (anonymous):

can u pls help me once more??

OpenStudy (anonymous):

Allrite just remember the formula for cos 2A=(1-tan^2A)/(1+tan^2A) Sec 2A=1/cos 2A So, Sec 2A=(1+tan^2A)/(1-tan^2A) Now Substitute for tan A

OpenStudy (anonymous):

k....another qstn ?

OpenStudy (anonymous):

if tan(πcos\[\theta\] = cot(πsin\[\theta\] , then cos(\[\theta\]- π/4) =

OpenStudy (anonymous):

Couldnot understand the question, pls type it properly so that I can help you

OpenStudy (anonymous):

problem with equation editor.. :( tan(πcos theta)= cot (π sin theta) , then cos(theta - π/4) =

OpenStudy (anonymous):

a. 2 b. 1/2√2

OpenStudy (anonymous):

Are you sure that the question is correct!

OpenStudy (anonymous):

yes...it θ

OpenStudy (anonymous):

i mean πcosθ

OpenStudy (anonymous):

It is b) but I am still working on it

OpenStudy (anonymous):

thanks 4 helping me...how did u get b?

myininaya (myininaya):

tan(a)=x/1=opp/adj sec(2a)=1/[cos(2a)]=1/[cos^2a-sin^2a]=1/[(cosa)^2-(sina)^2] hyp=sqrt{opp^2+adj^2}=sqrt{x^2+1} cosa=adj/hyp=1/[sqrt{x^2+1}] (cosa)^2=1/(x^2+1) sina=opp/hyp=x/[sqrt{x^2+1}] (sina)^2=x^2/[x^2+1] (cosa)^2-(sina)^2=(1-x^2)/(x^2+1) sec(2a)=(x^2+1)/(1-x^2)

OpenStudy (anonymous):

thank u .. :) can u pls help me with the other one too?

myininaya (myininaya):

ok im trying to get ready for something if i have time i will answer

OpenStudy (anonymous):

okie..=D